OpenStudy (anonymous):

The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (anonymous):

y = −2x − 2 y = -1/2x + 2 y = 1/2x − 2 y = − 2x + 2

OpenStudy (anonymous):

@MasterprojectHD @misty1212 @triciaal

OpenStudy (misty1212):

HI!! can you narrow it down to two?

OpenStudy (anonymous):

I will try

OpenStudy (misty1212):

let me ask it another way what is the slopeof \[y=\frac{1}{2}x+6\]?

OpenStudy (misty1212):

hint: the slope of \(y=mx+b\) is \(m\)

OpenStudy (anonymous):

1/2 is the slope? @misty1212

OpenStudy (misty1212):

yes and your parallel line will have the same slope, so you really only have one option

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

@misty1212